Award: 10.00 points Determine the accumulated and the discounted value of $1000 over 55 days at 7% using both ordinary and exact simple interest. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Ordinary interest Exact interest 2. Award: 10.00 points A loan of $2500 taken out on April 2 requires equal payments on May 25, July 20, September 10, and a final payment of $500 on October 15. If the focal date is October 15, what is the size of the equal payments at 9% simple interest? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Size of the equal payments 3. Award: 10.00 points Paul borrows $4000 at a 7.5% simple interest rate. He is to repay the loan by paying $1000 at the end of 3 months and two equal payments at the end of 6 months and 9 months. a) Determine the size of the equal payments using the end of 6 months as a focal date. (Do not round intermediate calculations. Round your answer to 2 decimal places.) Size of the equal payments b) Determine the size of the equal payments using the present time as a focal date. (Do not round intermediate calculations. Round your answer to 2 decimal places.]) Size of the equal payments
Answer #1

Answer1.

Accumulated value= 1000+1000×0.07×55/365=1010.54

Discounted value=(1000)/(1+0.07×55/365)=989 approx.

2.Total time period between april 2 to October 15=197 days

april= 29 days

may=31 days and so on

principle=2500

rate=9

total simple interest= 2500×0.09×197/365=121 approx

Total amount including interest to be paid= 2500+121=2621

equal installments= (2621-500( to be paid in oct))/3=707.

3.Borrowed amount=4000

rate=7.5

interest upto 3 month= 4000×7.5%×3/12=75

amount repaid after 3 month=1000

remaing amount 4000-1075=2925

.Amount to be repaid in equal installment in 6 month =

interest for 12 month = 2925*0.075*6/12=110 approx

amount repaid for 12 month = 2925+(110*2)=3145

equal installments of 6 month each= 3145/2=1572.

B)using present time as base equal time = 2925 as calculated above

interest for 15 month( 6 +9)= 2925*0.075*15/12=274

total amount=2925+274=3199

equal installments= 3199/2=1600